Four times the difference between a number and 5 is equal to the number increased by 4 What is the number?
Answer by 1335(25)

Four times the difference of a number and 5 means equates to: 4(x-5), with x being the unknown number.
You also know that this equation is equal to the number, x, increased by 4. So you now know that 4(x+5) is equal to (x+4). Now distribute the four in the first equation and get 4x-20. You also know that 4x-20 is also equal to x+4 or "4x-20=x+4". You can now simplify by subtracting an "x" from both sides, which will give you 3x-20=4. You can now add twenty to both sides and get 3x=24. Finally, divide both sides by 3 and you will get the answer: "x equals 8".